### 31 Mar 2026 v2

# Catalytic Quantum Error Correction: Theory, Efficient Catalyst Preparation,and Numerical Benchmarks

Abstract
We introduce Catalytic Quantum Error Correction (CQEC), a state recovery protocol exploiting catalytic covariant transformations [31]. CQEC recovers a known target state from noisy copies without an error magnitude threshold: recovery succeeds whenever the coherent modes satisfy $\mathcal{C}(\rho_{0})\subseteq\mathcal{C}(\rho_{\mathrm{noisy}})$, regardless of noise strength. The main practical bottleneck—catalyst preparation requiring $n^{*}\sim d^{4}e^{2\gamma}$ copies—is resolved by a three-stage pipeline combining CPMG dynamical decoupling, Clifford twirling, and the recursive swap test, achieving $F_{\mathrm{cat}}>0.96$ with only 8 copies ($10^{9}$-fold reduction). Numerical validation across four quantum algorithms (d=4d=46464), a cryptographic protocol, and three noise models confirms $F>0.999$ in the asymptotic limit across 200 configurations.
I Introduction
Quantum computers promise exponential speedups for problems in cryptography [29, 33], quantum simulation [6, 13], and machine learning [18, 4], but their practical realization is hindered by the fragility of quantum coherence under environmental decoherence [43, 28]. Quantum error correction (QEC) addresses this challenge by encoding logical information redundantly across physical qubits, enabling the detection and correction of errors below a code-specific threshold [32, 15]. However, the qubit overhead of conventional QEC is substantial: the Steane [[7,1,3]][\![7,1,3]\!] code requires 7 physical qubits per logical qubit [35], while surface codes at distance $d$ require $O(d^{2})$ qubits [20, 14], and correction fails entirely when error rates exceed the threshold $p_{\mathrm{th}}\approx 1\%$ [21].
A distinct approach to quantum state protection arises from the resource theory of quantum coherence [2, 36]. Coherence—the presence of off-diagonal elements in a state’s density matrix with respect to a preferred basis—is a quantifiable resource under the framework of covariant (energy-conserving) operations [25, 23]. Shiraishi and Takagi [31] recently established that the transformation rate between coherent states can diverge arbitrarily: given $n$ copies of a weakly coherent state $\rho$, one can produce $m\gg n$ copies of a target state $\rho^{\prime}$ via catalytic covariant operations, provided that the coherent modes of $\rho^{\prime}$ are contained within those of $\rho$. This result reveals that the output-to-input copy ratio $m/n$ can be made arbitrarily large (the asymptotic rate diverges), creating an infinitely sharp boundary between recoverable and irrecoverable quantum states—a feature with no analog in conventional QEC.
The role of coherence in quantum thermodynamics and resource theories has been extensively studied [23, 36, 25], with catalytic transformations appearing in the context of thermodynamic state conversion [2] and entanglement manipulation. The interplay between symmetry constraints and quantum error correction has been studied in the context of covariant codes [12] and approximate QEC [16]. Covariant codes [12] encode information in a way that respects a symmetry group, with an inherent trade-off between code accuracy and the symmetry charge variance (the Eastin–Knill theorem for exact codes [11]). CQEC differs fundamentally: it does not encode information redundantly but instead uses catalytic transformations to amplify coherence in the original Hilbert space. However, the specific application of catalytic coherence amplification to quantum error correction has not been explored. In particular, it remains unclear whether the infinite amplification rate translates into effective protection for realistic quantum algorithms operating under decoherence, what the finite-copy resource costs are, and how these costs compare with conventional QEC.
Moreover, even assuming the theoretical validity of catalytic recovery, the practical bottleneck is catalyst preparation. The constructive proof of Ref. [31] requires $n^{*}\sim d^{4}e^{2\gamma}$ noisy copies to prepare a catalyst with fidelity $F_{\mathrm{cat}}\geq 0.99$, where $d$ is the Hilbert space dimension and $\gamma$ is the dephasing strength. For a 6-qubit system ($d=64$) under moderate dephasing ($\gamma=2$), this translates to $n^{*}\approx 7.3\times 10^{10}$ copies—far beyond current experimental capabilities. Recent advances in quantum state purification [8, 41, 42] and dynamical decoupling [37, 26] offer paths to dramatically better scaling.
In this paper, we bridge these gaps by (i) introducing the Catalytic Quantum Error Correction (CQEC) protocol with explicit success conditions, (ii) developing a quantum circuit architecture based on energy-conserving gates, (iii) proposing four catalyst preparation strategies culminating in a DD+Twirl+Swap Test pipeline that achieves a $10^{9}$-fold copy-count reduction, and (iv) providing comprehensive numerical validation. Our central contributions are:
  1. 1.
    Protocol and circuit. We formulate the CQEC protocol with explicit success/failure conditions based on mode inclusion, implement it via energy-conserving gates in a minimal 4-qubit circuit, and provide finite-copy fidelity bounds.
  2. 2.
    Efficient catalyst preparation. We develop four catalyst preparation strategies culminating in the DD+Twirl+Swap Test pipeline, which achieves a $10^{9}$-fold reduction in copy count over distillation under strong dephasing.
  3. 3.
    Comprehensive benchmarks. We validate CQEC across four quantum algorithms (d=4d=46464), a cryptographic protocol, and three noise models (200 configurations), including comparison with conventional QEC codes, threshold verification across 10 orders of magnitude, and 100-cycle catalyst reuse.
While Theorems 1 and 2 are due to Shiraishi and Takagi [31], the application to quantum error correction, the circuit architecture, the catalyst preparation pipeline, and the comprehensive numerical validation are original to this work.
II Theoretical Framework
II.1 Unspeakable coherence and covariant operations
We work within the resource theory of unspeakable coherence [31, 24, 25, 23], where the free operations are covariant channels—completely positive trace-preserving maps $\Lambda$ satisfying
$\Lambda\circ\mathcal{U}_{t}=\mathcal{U}_{t}\circ\Lambda\quad\forall\,t\in\mathbb{R},$ (1)
where $\mathcal{U}_{t}(\rho)=e^{-iHt}\rho\,e^{iHt}$ is the time evolution generated by the system Hamiltonian $|H=\sum_{i}E_{i}\ket{i}\!\bra{i}$. Condition (1) enforces energy conservation: $U,H_{\mathrm{total}}]=0$ for any unitary $U$ implementing $\Lambda$ [25]. This is physically motivated by the Wigner–Araki–Yanase theorem, which constrains measurements and operations that do not commute with conserved quantities [39].
For a $d$-level system, the $\ell_{1}$-norm of coherence quantifies the total off-diagonal content:
$C_{\ell_{1}}(\rho)=\sum_{i\neq j}|\rho_{ij}|.$ (2)
The quantum Fisher information $\mathcal{F}(\rho,H)=2\sum_{i,j}\frac{(\lambda_{i}-\lambda_{j})^{2}}{\lambda_{i}+\lambda_{j}}|\!\braket{i|H|j}\!|^{2}$ [5] provides a complementary coherence monotone under covariant operations [31]. Unlike $C_{\ell_{1}}$, QFI is operationally meaningful for metrology and satisfies $\mathcal{F}(\rho_{\mathrm{noisy}})\leq e^{-2\gamma}\mathcal{F}(\rho_{0})$ under dephasing. We use QFI in Sec. VI.2 to quantify the coherence available for recovery. The two measures are complementary: $C_{\ell_{1}}$ counts total off-diagonal weight (relevant for mode coverage), while QFI weights contributions by their metrological utility. Under dephasing, both decay exponentially but at different rates: $C_{\ell_{1}}\sim e^{-\gamma}$ (slowest mode), $\mathcal{F}\sim e^{-2\gamma}$ (quadratic suppression), making QFI a more sensitive indicator of coherence loss.
II.2 Coherent modes and mode inclusion
The modes of asymmetry [23, 24] of a state $\rho$ are defined as
$\mathcal{D}(\rho)=\bigl\{\Delta_{ij}=E_{i}-E_{j}\mid\rho_{ij}\neq 0\bigr\}.$ (3)
The resonant coherent modes $\mathcal{C}(\rho)$ are the integer-linear span of $\mathcal{D}(\rho)$:
$\mathcal{C}(\rho)=\biggl\{\sum_{(i,j)}n_{ij}\,\Delta_{ij}\;\bigg|\;n_{ij}\in\mathbb{Z},\;\Delta_{ij}\in\mathcal{D}(\rho)\biggr\}.$ (4)
$\mathcal{C}(\rho)$ forms a subgroup of $(\mathbb{R},+)$ and determines the transformability of states under covariant operations [31].
Theorem 1 (Shiraishi–Takagi [31]). If $\mathcal{C}(\rho^{\prime})\subseteq\mathcal{C}(\rho)$ and $\rho$ is full rank, the asymptotic marginal transformation rate
$R(\rho\to\rho^{\prime})=\sup\biggl\{\frac{m}{n}\;\bigg|\;\rho^{\otimes n}\xrightarrow{\mathrm{cov}}\sigma,\;\mathrm{Tr}_{[m]\setminus\{k\}}[\sigma]\approx\rho^{\prime}\;\forall k\in[m]\biggr\}$
diverges: $R(\rho\to\rho^{\prime})=\infty$.
Theorem 2 (Shiraishi–Takagi [31]). For any $\rho,\rho^{\prime}$ with $\mathcal{C}(\rho^{\prime})\subseteq\mathcal{C}(\rho)$, there exists a catalyst $c$ and a covariant operation $\Lambda$ such that
$\mathrm{Tr}_{C}\bigl[\Lambda(\rho\otimes c)\bigr]=\rho^{\prime},\quad\mathrm{Tr}_{S}\bigl[\Lambda(\rho\otimes c)\bigr]=c.$ (5)
That is, $\rho\to\rho^{\prime}$ is achievable by correlated-catalytic covariant transformation.
Crucially, the mode inclusion condition is both necessary and sufficient: if $\mathcal{C}(\rho^{\prime})\not\subseteq\mathcal{C}(\rho)$, then $R(\rho\to\rho^{\prime})=0$ and no catalytic transformation can achieve the conversion [31]. This provides a sharp classification of state pairs into recoverable ($R=\infty$) and irrecoverable ($R=0$), with no intermediate regime.
We note that Theorem 2 establishes correlated catalysis: the output state $\tau=\Lambda(\rho\otimes c)$ satisfies $\mathrm{Tr}_{C}[\tau]=\rho^{\prime}$ and $\mathrm{Tr}_{S}[\tau]=c$, but $\tau\neq\rho^{\prime}\otimes c$ in general (the system and catalyst may be correlated in $\tau$). This is weaker than strict catalysis ($\tau=\rho^{\prime}\otimes c$), which imposes additional constraints [31]. For CQEC, correlated catalysis suffices: the catalyst is reusable because its reduced state is preserved, regardless of correlations with the recovered system. However, correlations between the catalyst and recovered system have operational consequences: (i) the system–catalyst correlations in $\tau$ mean that the recovered state $\rho^{\prime}=\mathrm{Tr}_{C}[\tau]$ may differ from the state obtained by measuring the catalyst and post-selecting, and (ii) over multiple reuse cycles, correlations could accumulate in the joint state, potentially degrading effective catalyst quality. In our 100-cycle durability test (Sec. VI.6), the catalyst deviation remains below $10^{-12}$ per cycle, indicating that correlation accumulation is negligible at this system size.
II.3 CQEC protocol
The CQEC protocol applies Theorem 2 to quantum state recovery. We use the term “recovery” rather than “correction” to emphasize the key operational distinction: CQEC recovers a known target state from noisy copies, whereas QEC corrects errors on an unknown encoded state. The requirement that $\rho_{0}$ be known does not reduce CQEC to trivial state re-preparation: the catalyst enables recovery from noisy copies without re-executing the (potentially expensive) state preparation circuit, provided sufficient copies are available. CQEC does not violate the no-cloning theorem [40] for three reasons: (i) the target state $\rho_{0}$ is known, so no unknown-state information is being duplicated; (ii) the $n$ input copies are consumed (not preserved) by the protocol; and (iii) the covariant channel $\Lambda$ is state-specific and cannot be applied to arbitrary unknown states.
Given an ideal state $\rho_{0}$ corrupted by decoherence to $\rho_{\mathrm{noisy}}$, the protocol proceeds as follows.
Step 0 (Catalyst construction). Construct a full-rank catalyst $c$ with $\mathcal{D}(c)\supseteq\mathcal{D}(\rho_{0})$, following the construction of Proposition S.23 in Ref. [31]:
[…]
where $n_{\mathrm{cat}}$ is the number of catalytic branches, $\tau_{n_{\mathrm{cat}}-k}=\mathrm{Tr}_{1,\ldots,k}[\sigma_{n_{\mathrm{cat}}}]$ is the reduced state, and $R$ is a register system of dimension $n_{\mathrm{cat}}$. The apparent circularity is resolved by the constructive proof of Ref. [31]: two-level catalysts $c_{i}$ are first constructed from $\mu_{i}$ copies via non-catalytic covariant operations, then composed into the full catalyst via Theorem S.10. The catalyst dimension is $\dim(c)=n_{\mathrm{cat}}\cdot d^{n_{\mathrm{cat}}-1}\cdot n_{\mathrm{cat}}$, which grows exponentially. In our variational implementation (Sec. III.3), we use a simplified catalyst of dimension $d_{C}=2^{n_{C}}$ with $n_{C}=n_{S}$ qubits. This gap is fundamental: the variational circuit searches for a high-fidelity covariant transformation within a restricted ansatz, validating theoretical predictions but not implementing the full constructive proof. Efficient catalyst preparation strategies that dramatically reduce the required resources are developed in Sec. IV.
Step 1 (Mode verification). Compute $\mathcal{D}(\rho_{\mathrm{noisy}})$ and verify $\mathcal{C}(\rho_{0})\subseteq\mathcal{C}(\rho_{\mathrm{noisy}})$. In our numerical implementation, a mode $\Delta_{ij}$ is considered present if $|\rho_{ij}|>\epsilon_{\mathrm{mode}}$ with threshold $\epsilon_{\mathrm{mode}}=10^{-14}$, chosen to be two orders of magnitude above 64-bit machine epsilon. For partial dephasing $\rho_{ij}\to\rho_{ij}e^{-\gamma|\Delta_{ij}|}$ with $\gamma<\infty$, all modes survive: $\mathcal{D}(\rho_{\mathrm{noisy}})=\mathcal{D}(\rho_{0})$. For complete dephasing ($\gamma=\infty$), $\mathcal{D}(\rho_{\mathrm{noisy}})=\emptyset$ and recovery is impossible.
Step 2 (Catalytic recovery). Apply the covariant operation $\Lambda$ from Theorem 2 to execute
$\Lambda(\rho_{\mathrm{noisy}}\otimes c)=\tau,\quad\text{with}\quad\mathrm{Tr}_{C}[\tau]\approx\rho_{0},\quad\mathrm{Tr}_{S}[\tau]=c.$ (7)
Step 3 (Catalyst reuse). Since $\mathrm{Tr}_{S}[\tau]=c$, the catalyst is unchanged and available for the next recovery cycle, enabling infinite reuse.
Corollary (CQEC success condition). CQEC recovers $\rho_{0}$ from $\rho_{\mathrm{noisy}}$ if and only if $\mathcal{C}(\rho_{0})\subseteq\mathcal{C}(\rho_{\mathrm{noisy}})$ and $\rho_{\mathrm{noisy}}$ is full rank. Here the full-rank condition on $\rho_{\mathrm{noisy}}$ is required by Theorem 1 for $R=\infty$; Theorem 2 (catalytic single-copy conversion) does not require full rank but produces only correlated-catalytic output. Pure algorithmic states $|\ket{\psi}\!\bra{\psi}$ have rank 1, but after any of our noise channels with finite parameters, $\rho_{\mathrm{noisy}}$ acquires full rank. The threshold is infinitely sharp: for any family parametrized by residual coherence $\varepsilon$ with $\mathcal{C}(\rho_{0})\subseteq\mathcal{C}(\rho_{\mathrm{noisy}}(\varepsilon))$ for $\varepsilon>0$ and $\mathcal{C}(\rho_{0})\not\subseteq\mathcal{C}(\rho_{\mathrm{noisy}}(0))$, we have $F\to 1$ for $\varepsilon>0$ and $F=1/d$ for $\varepsilon=0$.
II.4 Copy scaling
From the proof of Theorem S.10 in Ref. [31], the full protocol requires:
  • Input copies: $n=\mu\cdot N^{k}$, where $\mu=\sum_{i=1}^{N}\mu_{i}$ is the total catalyst preparation cost and $N$ is the number of two-level catalysts.
  • Output copies: $m=(k+1)\cdot N^{k}$.
  • Rate: $R=m/n=(k+1)/\mu\to\infty$ as $k\to\infty$.
Each $\mu_{i}$ depends on the weakest coherent mode of the source state [31]:
$\mu_{i}\sim\frac{1}{|\rho_{ij,\min}|^{2}},$ (8)
where $\rho_{ij,\min}$ is the smallest nonzero off-diagonal element. For dephased states with decay factor $e^{-\gamma}$, $\mu_{i}\sim e^{2\gamma}/|\rho_{ij}^{(0)}|^{2}$, showing exponential growth in dephasing strength.
III Quantum Circuit Architecture
III.1 Energy-conserving gate
The fundamental gate for CQEC is the energy-conserving (EC) rotation [30, 31]:
$U_{\mathrm{EC}}(\theta)=\begin{pmatrix}1&0&0&0\\ 0&\cos\theta&-i\sin\theta&0\\ 0&-i\sin\theta&\cos\theta&0\\ 0&0&0&1\end{pmatrix}$, (9)
[…]
III.2 Three-layer circuit structure
The CQEC circuit operates on four registers: system ($S$, $n_{S}$ qubits), catalyst ($C$, $n_{C}$ qubits), and two ancillae ($A_{0},A_{1}$). The circuit applies EC gates in three layers:
Layer 1 (System–Catalyst, $n_{S}\times n_{C}$ gates): Transfers coherence from the system to the catalyst:
$U_{\mathrm{L1}}=\prod_{s\in S}\prod_{c\in C}U_{\mathrm{EC}}(\theta_{sc}).$ (10)
Layer 2 (Catalyst–Ancilla, $n_{C}\times n_{A}$ gates): Distributes coherence from the catalyst to ancillae:
$U_{\mathrm{L2}}=\prod_{c\in C}\prod_{a\in A}U_{\mathrm{EC}}(\theta_{ca}).$ (11)
Layer 3 (System–Ancilla, $n_{S}\times n_{A}$ gates): Direct coherence amplification between system and ancillae:
$U_{\mathrm{L3}}=\prod_{s\in S}\prod_{a\in A}U_{\mathrm{EC}}(\theta_{sa}).$ (12)
[…]
III.3 Minimal 4-qubit implementation
For $n_{S}=1$, $n_{C}=1$, $n_{A}=2$, the circuit uses 5 EC gates with parameters $\theta_{0}$ through $\theta_{4}$ (Fig. 1): Layer 1 applies 1 gate ($\theta_{0}$) for coherence bridging; Layer 2 applies 2 gates ($\theta_{1},\theta_{2}$) for coherence distribution; Layer 3 applies 2 gates ($\theta_{3},\theta_{4}$) for direct amplification.
[…]
III.4 Parameter optimization
The gate parameters $\boldsymbol{\theta}=(\theta_{0},\ldots,\theta_{4})$ are optimized to maximize a combined objective:
$\mathcal{L}(\boldsymbol{\theta})=0.7\cdot F(\rho_{S}^{\mathrm{out}},\rho_{0})+0.3\cdot F(\rho_{C}^{\mathrm{out}},c),$ (13)
where $F(\rho,\sigma)=\bigl(\mathrm{Tr}\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\bigr)^{2}$ is the Uhlmann fidelity [19], $\rho_{S}^{\mathrm{out}}=\mathrm{Tr}_{CA}[U_{\mathrm{CQEC}}(\rho_{\mathrm{in}})U_{\mathrm{CQEC}}^{\dagger}]$ is the recovered system state, and $\rho_{C}^{\mathrm{out}}=\mathrm{Tr}_{SA}[\cdot]$ is the catalyst after the operation. The weighting $\alpha=0.7$ reflects the asymmetry of the catalytic constraint: the catalyst must be exactly preserved (a hard constraint), while recovery fidelity should be maximized (the objective). The $\alpha=0.7$ value places sufficient weight on system recovery while keeping the catalyst penalty high enough to prevent drift. Scanning $\alpha\in[0.5,0.9]$ confirms $\alpha\in[0.6,0.8]$ maintains both $F_{\mathrm{after}}>0.999$ and $F_{C}>0.998$ across all tested combinations; outside this range, either the recovered state degrades ($\alpha<0.6$) or the catalyst drifts ($\alpha>0.8$).
Optimization uses a two-stage gradient-free approach: (i) Latin hypercube sampling of 100 initial parameter vectors $\boldsymbol{\theta}\in[0,2\pi)^{5}$, followed by (ii) Nelder–Mead simplex refinement [27] of the best 5 candidates. The parameters are optimized for each specific input state: CQEC in its current variational form is a state-specific decoder. Convergence is assessed via 50 independent random initializations; in all cases, the final objective $\mathcal{L}>0.98$ is achieved within 120 iterations with variance $10^{-6}$.
Relation to the formal construction. The 5-gate circuit is a variational ansatz designed to approximate the covariant channel $\Lambda$ of Theorem 2, not a direct implementation of the constructive proof. The EC gate structure ensures covariance by construction, while the optimized parameters are tuned for each specific input state. Layered EC gates do not form a universal gate set for covariant unitaries; however, universality is not required—we only need sufficient expressibility for the specific recovery map $\Lambda$. The dimension of the covariant unitary group for $n_{q}$ qubits is $\sum_{k}\binom{n_{q}}{k}^{2}-1$ [30]; our ansatz explores a $5$-dimensional submanifold, yet achieves $F>0.999$ for $d\leq 64$, implying the target $\Lambda$ lies near this submanifold. For states outside the ansatz’s reachable set, increasing the gate count to $\geq 15$ extends coverage (Sec. III.4).
Approximation quality vs. dimension. The fidelity gap grows with $d$: $1-F<10^{-8}$ for $d=4$, $1-F<10^{-4}$ for $d=8$, $1-F\approx 10^{-4}$ for $d=16$, and $1-F\approx 2\times 10^{-4}$ for $d=64$. Increasing to 10–15 gates recovers $10^{-6}$ for all $d$. The covariant unitary group for $n=8$ qubits has $\sum_{k}\binom{n}{k}^{2}-1=12\,869$ free parameters [30]; our 5–15 parameter ansatz explores a small submanifold, with high fidelity implying the target $\Lambda$ lies near this submanifold.
IV Efficient Catalyst Preparation
IV.1 Resource bottleneck
The finite-copy fidelity bound from Ref. [31] gives
$1-F(\rho_{S}^{\mathrm{out}},\rho_{0})\leq\frac{C^{2}}{4n}+\frac{C}{\sqrt{n}},$ (14)
where
$C\leq\frac{d\cdot C_{\ell_{1}}(\rho_{0})}{\displaystyle\min_{(i,j)\in\mathcal{D}(\rho)}|\rho_{ij}|}.$ (15)
For dephased states, $\min|\rho_{ij}|$ decays as $e^{-\gamma(d-1)}$, so $C$ grows exponentially with both $d$ and $\gamma$. Table 1 shows the prohibitive cost for realistic parameters.
Table 1: Copy count $n^{*}$ for $F_{\mathrm{cat}}\geq 0.99$ via Shiraishi–Takagi distillation ($\gamma=2$).
Algorithm $d$ $C$ $n^{*}$
QKAN 4 8.5 $7.3\times 10^{5}$
qDRIFT 8 42 $1.8\times 10^{7}$
CF-QPE 16 170 $2.9\times 10^{8}$
Regev 64 2700 $7.3\times 10^{10}$
IV.2 Variational catalyst preparation
We parameterize the catalyst state via a product of energy-conserving two-level rotations:
$U(\vec{\theta})=\prod_{\ell=1}^{L}\prod_{i<j}G_{ij}(\theta_{ij}^{(\ell)},\phi_{ij}^{(\ell)}),$ (16)
where $G_{ij}(\theta,\phi)$ acts on the $(i,j)$ subspace as
$G_{ij}=\begin{pmatrix}\cos\theta&-e^{i\phi}\sin\theta\\ e^{-i\phi}\sin\theta&\cos\theta\end{pmatrix}$ (17)
[…]
The optimization minimizes
$\mathcal{L}(\vec{\theta})=-w_{1}\frac{C_{\ell_{1}}(\rho_{\mathrm{cat}})}{d-1}+w_{2}\frac{|\mathcal{M}_{\mathrm{missing}}|}{|\mathcal{M}_{\mathrm{target}}|}-w_{3}\log\rho_{\mathrm{min}},$ (18)
where $\mathcal{M}_{\mathrm{missing}}$ is the set of target modes absent from $\rho_{\mathrm{cat}}$, $\rho_{\mathrm{min}}=\min_{i}(\rho_{\mathrm{cat}})_{ii}$, and $(w_{1},w_{2},w_{3})=(1,10,5)$. The weight hierarchy $w_{2}\gg w_{3}>w_{1}$ reflects priority: mode coverage >> population balance >> raw coherence. A sensitivity analysis over 20 weight combinations ($w_{1}\in[0.5,2]$, $w_{2}\in[5,20]$, $w_{3}\in[2,10]$) shows that $C_{\ell_{1}}/(d-1)>0.95$ and 100% mode coverage are achieved for all combinations with $w_{2}\geq 5$; the population balance ($\rho_{\mathrm{min}}$) varies by at most $2\times$, affecting downstream recovery by $\Delta F_{\mathrm{rec}}<0.03$. L-BFGS-B optimization with 5 random restarts converges in <5<5 seconds for $d\leq 16$.
Table 2: Variational catalyst preparation. Zero copies consumed.
$d$ $N_{\theta}$ $C_{\ell_{1}}$ Mode cov. $F_{\mathrm{rec}}$ (deph) $F_{\mathrm{rec}}$ (depol)
4 36 3.00 100% 0.83 0.87
8 168 6.99 100% 0.73 0.78
16 720 14.5 100% 0.65 0.71
The variational approach achieves maximal $\ell_{1}$-coherence and 100% mode coverage for $d\leq 16$, consuming no noisy copies—suitable for NISQ devices where copies are expensive. However, the $N_{\theta}\sim O(d^{2})$ parameter count makes optimization intractable for $d\geq 32$.
IV.3 Recursive swap test purification
The swap test [8, 1] on two copies $\rho$ and $\sigma$ projects onto the symmetric subspace $\mathrm{Sym}^{2}(\mathbb{C}^{d})$ via the projector $\Pi_{\mathrm{sym}}=(I+\mathrm{SWAP})/2$:
$\omega=\frac{\rho+\sigma+\rho\sigma+\sigma\rho}{2(1+\mathrm{Tr}(\rho\sigma))}.$ (19)
In the recursive protocol, both inputs are identically prepared ($\sigma=\rho$). Substituting into Eq. (19): $\omega=(\rho+\rho+\rho^{2}+\rho^{2})/(2(1+\mathrm{Tr}(\rho^{2})))=(\rho+\rho^{2})/(1+\mathrm{Tr}(\rho^{2}))$. For depolarized states $\rho(\delta)=(1-\delta)\psi+\delta I/d$, recursive application yields error $\delta_{k}\sim\delta^{2^{k}}$—doubly exponential convergence [8].
We note the structural similarity to entanglement purification [3], which also uses bilateral operations and post-selection on two copies; however, coherence purification operates on a single system and the relevant resource is $\ell_{1}$-coherence rather than entanglement.
Under depolarizing noise ($p=0.3$), the standard recursive swap test achieves rapid convergence (Table 3).
Table 3: Standard recursive swap test under depolarizing noise ($p=0.3$). $F_{\mathrm{cat}}$ values shown.
$d$ $n=8$ $n=32$ $n=64$ $n^{*}$ (distillation)
4 0.949 0.986 0.993 $7.3\times 10^{5}$
8 0.940 0.983 0.991 $1.8\times 10^{7}$
16 0.935 0.982 $2.9\times 10^{8}$
64 0.932 $7.3\times 10^{10}$
Under energy-dependent dephasing ($\gamma=2$), the standard swap test fails severely (Table 4).
Table 4: Standard recursive swap test under dephasing ($\gamma=2$). $F_{\mathrm{cat}}$ at maximum tested copy count $n_{\mathrm{max}}$.
Algorithm $d$ $F_{\mathrm{cat}}$ ($n_{\mathrm{max}}$) $C_{\ell_{1}}$
QKAN 4 0.438 (64) 1.52
qDRIFT 8 0.195 (64) 2.94
CF-QPE 16 0.088 (32) 4.54
Regev 64 0.021 (8) 12.5
[…]
IV.4 Dynamical decoupling and Clifford twirling
The root cause of purification failure under dephasing is that dephasing is anisotropic: it suppresses coherences $\rho_{ij}$ proportionally to $e^{-\gamma|E_{i}-E_{j}|}$, creating an exponentially non-uniform error profile that the swap test’s symmetric projection cannot efficiently correct. We introduce a three-stage pipeline that overcomes this limitation.
CPMG dynamical decoupling. The CPMG sequence [26, 37] applies $N$ equally spaced $\pi$-pulses during the dephasing interval, reducing the effective dephasing to
$\gamma_{\mathrm{eff}}=\frac{\gamma}{N+1},$ (20)
[…]
Table 5: $D$ sweep for qDRIFT ($d=8$, $n=8$ copies, $\gamma=2$). $\gamma_{\mathrm{eff}}$ is the effective dephasing after $D$, $p_{\mathrm{eff}}$ is the depolarizing parameter after twirling, and $F_{\mathrm{cat}}$ is the catalyst fidelity after the full pipeline.
DD config $\gamma_{\mathrm{eff}}$ $p_{\mathrm{eff}}$ $F_{\mathrm{cat}}$
No DD 2.000 0.961 0.176
CPMG-2 0.667 0.541 0.800
CPMG-4 0.400 0.366 0.914
CPMG-8 0.222 0.221 0.963
CPMG-16 0.118 0.123 0.983
Even moderate DD pulse counts produce dramatic improvements: CPMG-2 lifts $F_{\mathrm{cat}}$ from 0.176 to 0.800, and CPMG-8 achieves $F_{\mathrm{cat}}=0.963$ (Fig. 3). The same trend holds across all tested dimensions: at CPMG-8, $F_{\mathrm{cat}}=0.964$ ($d=4$), $0.963$ ($d=8$, 1616), and $0.963$ ($d=64$), confirming dimension independence (Table 6).
Clifford twirling. After DD reduces $\gamma$ to $\gamma_{\mathrm{eff}}$, the residual noise is still anisotropic dephasing. Clifford twirling [10, 38] converts any noise channel into a depolarizing channel by averaging over random Clifford gates (a technique related to randomized benchmarking [10] and randomized compiling [38]):
$\overline{\mathcal{E}}_{\mathrm{twirl}}(\rho)=(1-p_{\mathrm{eff}})\rho+p_{\mathrm{eff}}\,\frac{I}{d},$ (21)
where
$p_{\mathrm{eff}}=1-\frac{d\,\bar{F}_{\mathrm{avg}}-1}{d-1},$ (22)
and $\bar{F}_{\mathrm{avg}}=d^{-2}\sum_{i,j}e^{-\gamma_{\mathrm{eff}}|E_{i}-E_{j}|}$ for dephasing with $H_{S}=\mathrm{diag}(0,1,\ldots,d-1)$. The twirling converts the exponentially non-uniform dephasing profile into a single depolarizing parameter $p_{\mathrm{eff}}$, which the swap test handles efficiently. In practice, full Clifford twirling requires averaging over all $|\mathrm{Cl}(d)|$ group elements; we approximate with $K$ random samples, incurring a sampling error $\epsilon_{\mathrm{twirl}}\sim d/\sqrt{K}$ [38]. For $d=64$ and $K=200$ random Cliffords per copy, $\epsilon_{\mathrm{twirl}}\approx 0.05$, which is subdominant to the residual depolarizing parameter $p_{\mathrm{eff}}\geq 0.22$. In our simulations, we use the exact twirled channel (the analytical average) to isolate the pipeline’s asymptotic performance from sampling artifacts.
IV.5 Combined DD+Twirl+Swap Test pipeline
The full pipeline operates in three stages:
  1. 1.
    DD stage: Apply CPMG-$N$ to each noisy copy, reducing $\gamma\to\gamma_{\mathrm{eff}}=\gamma/(N+1)$.
  2. 2.
    Twirl stage: Apply Clifford twirling to each DD-protected copy, converting dephasing($\gamma_{\mathrm{eff}}$) $\to$ depolarizing($p_{\mathrm{eff}}$).
  3. 3.
    Purification stage: Apply the standard recursive swap test, exploiting doubly exponential convergence under depolarizing noise.
The physical interpretation is: DD suppresses the noise magnitude, twirling isotropizes the residual noise geometry, and the swap test purifies the isotropic noise efficiently. Each stage addresses a distinct obstacle—magnitude, anisotropy, and mixedness—and the composition is more effective than any single technique.
Table 6 presents the main results with CPMG-8 and $n=8$ copies.
Table 6: DD+Twirl+Swap Test pipeline results (CPMG-8, $\gamma_{\mathrm{eff}}=0.222$, $n=8$ copies, original $\gamma=2$). “Raw $F_{\mathrm{cat}}$” is without DD or twirling.
Algorithm $d$ Raw $F_{\mathrm{cat}}$ DD+Twirl $F_{\mathrm{cat}}$ DD+Twirl $F_{\mathrm{rec}}$
QKAN 4 0.351 0.964 0.811
qDRIFT 8 0.173 0.963 0.900
CF-QPE 16 0.085 0.963 0.649
Regev 64 0.021 0.963 0.636
The pipeline achieves $F_{\mathrm{cat}}>0.96$ uniformly across all dimensions, in stark contrast to the raw swap test ($F_{\mathrm{cat}}=0.02$$0.44$), DD alone (dimension-dependent), or twirling alone (which fails because twirling without DD converts strong dephasing into strong depolarizing noise). An ablation analysis confirms that each stage is necessary: DD alone reduces $\gamma$ but leaves anisotropic noise ($F_{\mathrm{cat}}\approx 0.6$$0.8$, dimension-dependent); twirling alone converts strong dephasing into strong depolarizing ($p_{\mathrm{eff}}\approx 0.96$ for $\gamma=2$), which the swap test cannot efficiently purify; DD+Twirl without the swap test produces $F_{\mathrm{cat}}\approx 0.85$ (the twirled state’s overlap with the target); and the full pipeline achieves $F_{\mathrm{cat}}>0.96$ uniformly (Figs. 2 and 4).
The total resource cost is: $N$ DD pulses per copy (CPMG-8: 8 $\pi$-pulses per copy), $O(n_{q}^{2})$ Clifford gates per copy for twirling, and $n=2^{k}$ copies for the swap test ($k=3$, $n=8$). For $d=64$ ($n_{q}=6$), the pipeline requires 8 copies, 64 DD pulses, and $\sim\!288$ Clifford gates, compared to $7.3\times 10^{10}$ copies for distillation—a reduction factor of $7.3\times 10^{10}/8\approx 10^{10}$. We conservatively quote $10^{9}$-fold” to account for the additional gate overhead of DD and twirling, which effectively multiplies each copy’s preparation cost by a factor of $\sim\!10$. The copy counts required for $F_{\mathrm{cat}}\geq 0.99$ under the pipeline are $n=32$ for all tested $d\leq 16$, representing a 4–8 orders of magnitude reduction over distillation.
Figure 2: Pipeline comparison under dephasing $\gamma=2$ with $n=8$ copies (CPMG-8). Raw purification (red) fails for $d\geq 8$. DD alone (orange) improves but is dimension-dependent. Twirl alone (purple) fails. DD+Twirl (green) achieves $F_{\mathrm{cat}}>0.96$ uniformly.
Figure 3: Catalyst fidelity vs. CPMG pulse count for qDRIFT ($d=8$, $n=8$ copies, $\gamma=2$). DD reduces effective dephasing, twirling isotropizes, and the swap test purifies efficiently.
Figure 4: Summary of all catalyst preparation strategies under dephasing $\gamma=2$: variational (0 copies), standard swap test, covariant swap test, and DD+Twirl+Swap Test pipeline (CPMG-8, 8 copies). The DD+Twirl pipeline dramatically outperforms all other methods.
V Benchmark Algorithms and Decoherence Models
V.1 Quantum algorithms
We test CQEC on four algorithms chosen to span the major application domains and cover a range of Hilbert space dimensions (d=4d=46464), coherence structures, and algorithmic noise profiles.
1. qDRIFT [6, 7] simulates Hamiltonian dynamics $e^{-iHt}$ for the 3-qubit Heisenberg model $H=J\sum_{\langle i,j\rangle}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y}+\sigma_{i}^{z}\sigma_{j}^{z})+h\sum_{i}\sigma_{i}^{z}$ with $J=1.0$, $h=0.5$, $t=1.0$. The qDRIFT approximation uses 80 random product formula gates with probabilistic sampling [6]. Dimension $d=8$ (3 qubits). Since different random seeds produce different gate sequences, $F_{\mathrm{before}}$ varies across seeds (std $\approx 0.17$ for dephasing $\gamma=2$, reflecting the randomized product formula’s state-dependent approximation error). Crucially, the post-correction fidelity $F_{\mathrm{after}}$ shows negligible seed dependence (std $10^{-4}$), confirming that CQEC is robust to the input state’s seed-dependent structure.
2. QKAN [18] implements a quantum Kolmogorov–Arnold network layer encoding the first four Chebyshev polynomials $T_{n}(x)$ at $x=0.5$: the ideal state has amplitudes proportional to $(T_{0},T_{1},T_{2},T_{3})=(1.0,0.5,-0.5,-1.0)$ (normalized), while the algorithmic output truncates at degree 2. The CQEC target is the algorithmic output (not the ideal state), so recovery corrects only decoherence. Dimension $d=4$ (2 qubits).
3. Control-free QPE [9] estimates eigenvalues of a Fermi–Hubbard-type Hamiltonian using vectorial phase retrieval without controlled unitaries. The protocol generates time series $f_{j}=\braket{\psi|e^{-iHj\Delta t}|\psi}$ and encodes the resulting spectrum as a 16-dimensional quantum state. Dimension $d=16$ (4 qubits).
4. Regev factoring [29] factors $N=15$ using discrete Gaussian states with $D=8$ grid points per dimension, modular exponentiation, and quantum Fourier transform for LLL reduction [22]. Dimension $d=64$ (6 qubits).
V.2 Decoherence models
Three noise channels are applied to each algorithm’s output state:
Partial dephasing. $\mathcal{E}_{\mathrm{deph}}(\rho)_{ij}=\rho_{ij}\cdot e^{-\gamma|E_{i}-E_{j}|}$, with $\gamma=2.0$ (strong dephasing). This preserves all coherent modes for $\gamma<\infty$: $\mathcal{D}(\rho_{\mathrm{noisy}})=\mathcal{D}(\rho)$. For dephased states, the pre-correction fidelity with the ideal state is
$F_{\mathrm{before}}(\gamma)$ $=\mathrm{Tr}[\rho_{\mathrm{diag}}^{2}]+e^{-\gamma}\bigl(1-\mathrm{Tr}[\rho_{\mathrm{diag}}^{2}]\bigr),$ (23)
which approaches $1/d$ as $\gamma\to\infty$ for the maximally coherent state.
Depolarizing. $\mathcal{E}_{\mathrm{depol}}(\rho)=(1-p)\rho+p\cdot I/d$, with $p=0.3$. Since $I/d$ has zero off-diagonal elements, all coherent modes are preserved for $p<1$.
[…]
V.3 TTN cryptographic protocol
To demonstrate CQEC’s applicability beyond algorithm protection, we test it on quantum states generated by a tree tensor network (TTN) cryptographic protocol [17, 34]. We simulate a 3-qubit reduced model with layered RY–CNOT–RZ structure, tested across 6 plaintext lengths ($N_{c}=5,10,15,20,25,30$) and 7 noise levels from ideal to extreme.
V.4 Conventional QEC baselines
For comparison, we compute the logical fidelity of:
  • Steane [[7,1,3]][\![7,1,3]\!] code [35]: corrects any single-qubit error ($t=1$). $F_{L}=\sum_{k=0}^{t}\binom{7}{k}p_{\mathrm{eff}}^{k}(1-p_{\mathrm{eff}})^{7-k}$.
  • Surface code [20, 14] at distances $d=3$ and $d=5$: $p_{L}\sim(p/p_{\mathrm{th}})^{(d+1)/2}$ below $p_{\mathrm{th}}\approx 0.01$.
The effective per-qubit error rate is $p_{\mathrm{eff}}=1-(1-p)^{1/n_{q}}$. These are analytical models; both sides of the comparison represent idealized performance. The comparison illustrates qualitative differences: threshold-dependent degradation vs. threshold-free recovery. We emphasize that QEC codes correct arbitrary errors up to weight $t$ on unknown encoded states, while CQEC recovers known states from arbitrary noise strength—these are complementary capabilities, not competing ones.
VI Results
VI.1 CQEC asymptotic benchmark
We present results in two regimes: the asymptotic limit ($n\to\infty$, Secs. VI.1VI.6), which validates the existence and robustness of catalytic recovery, and the finite-copy regime (Secs. VI.7VI.9), which quantifies the practical resource cost. We caution that the asymptotic results represent an upper bound on achievable performance; Table 9 gives the copy counts needed to approach these fidelities, and the DD+Twirl pipeline (Sec. VI.8) dramatically reduces these costs. The two regimes should be read together: Sec. VI.1 answers “what is possible?” and Secs. VI.7VI.9 answer “at what cost?”
Table 7 presents CQEC recovery results in the asymptotic limit from 10 independent trials. Regev ($d=64$) achieves slightly lower $F_{\mathrm{after}}$ compared to smaller systems, reflecting the reduced expressibility of the 5-parameter ansatz in the larger Hilbert space.
Table 7: CQEC recovery fidelity (mean $\pm$ std, 10 seeds). Success rate is 100% for all entries. Regev results (below rule) serve as a high-dimensional stress test; CQEC is inapplicable to factoring in practice due to the target-state requirement (see Sec. VII.3).
Algorithm Noise model $F_{\mathrm{before}}$ $F_{\mathrm{after}}$
qDRIFT (3 qb) Dephasing ($\gamma=2$) $0.701\pm 0.171$ $0.9999\pm 0.0000$
qDRIFT (3 qb) Depolarizing ($p=0.3$) $0.528\pm 0.119$ $0.9997\pm 0.0001$
qDRIFT (3 qb) Combined $0.615\pm 0.145$ $0.9999\pm 0.0001$
QKAN (2 qb) Dephasing ($\gamma=2$) $0.341\pm 0.000$ $1.0000\pm 0.0000$
QKAN (2 qb) Depolarizing ($p=0.3$) $0.495\pm 0.000$ $1.0000\pm 0.0000$
QKAN (2 qb) Combined $0.386\pm 0.000$ $1.0000\pm 0.0000$
CF-QPE (4 qb) Dephasing ($\gamma=2$) $0.306\pm 0.000$ $1.0000\pm 0.0000$
CF-QPE (4 qb) Depolarizing ($p=0.3$) $0.718\pm 0.000$ $1.0000\pm 0.0000$
CF-QPE (4 qb) Combined $0.427\pm 0.000$ $1.0000\pm 0.0000$
TTN-Crypto (3 qb) Dephasing ($\gamma=2$) $0.388\pm 0.067$ $1.0000\pm 0.0000$
TTN-Crypto (3 qb) Depolarizing ($p=0.3$) $0.734\pm 0.002$ $1.0000\pm 0.0000$
TTN-Crypto (3 qb) Combined $0.487\pm 0.042$ $1.0000\pm 0.0000$
Regev (6 qb) Dephasing ($\gamma=2$) $0.066\pm 0.000$ $0.9998\pm 0.0000$
Regev (6 qb) Depolarizing ($p=0.3$) $0.505\pm 0.000$ $0.9996\pm 0.0001$
Regev (6 qb) Combined $0.271\pm 0.000$ $0.9997\pm 0.0001$
VI.2 Sharp threshold
Figure 5 demonstrates the infinitely sharp zero/nonzero threshold predicted by Theorem 1. For a 2-qubit ($d=4$) maximally coherent target state:
  • $\varepsilon=0$: $F_{\mathrm{after}}=0.250=1/d$. Recovery fails.
  • $\varepsilon=10^{-10}$: $F_{\mathrm{after}}=1.0000$. Full recovery despite $C_{\ell_{1}}=1.2\times 10^{-9}$. At this coherence level, off-diagonal matrix elements are $O(10^{-10})$, well above 64-bit machine epsilon ($\approx 2.2\times 10^{-16}$) and the mode detection threshold $\epsilon_{\mathrm{mode}}=10^{-14}$ (Sec. II.3).
  • All $\varepsilon>0$ (30 values on $10^{-10},0.32]$): $F_{\mathrm{after}}=1.000$.
The QFI confirms quantitatively: at $\varepsilon=10^{-10}$, $\mathcal{F}(\rho_{\mathrm{noisy}},H)=3.2\times 10^{-19}$ yet recovery succeeds, while at $\varepsilon=0$, $\mathcal{F}=0$ and recovery fails. This demonstrates that the threshold is controlled by the support of the coherence, not its magnitude.
We also tested a selective dephasing channel that destroys the $\Delta_{01}$ mode while preserving all others: CQEC achieves only $F_{\mathrm{after}}=0.72$ (vs. $F_{\mathrm{after}}=1.0$ with all modes present). The intermediate value $F=0.72>1/d=0.25$ reflects partial recovery: the optimizer finds a covariant map that reconstructs the state using the surviving modes, but cannot restore the lost $\Delta_{01}$ component. This demonstrates that mode inclusion violations produce graceful degradation rather than catastrophic failure.
Figure 5: Sharp threshold of CQEC. Post-correction fidelity vs. residual coherence $\varepsilon$ (log scale). Green circles: successful recovery ($\varepsilon>0$). Red cross: failed recovery ($\varepsilon=0$). Dashed line: $F=1/d=0.25$.
VI.3 Noise strength sweep
Figure 6 presents CQEC recovery fidelity vs. noise strength. For dephasing with $\gamma\in[0.1,5.0]$ (20 points) and depolarizing with $p\in[0.01,0.95]$ (20 points), CQEC maintains $F_{\mathrm{after}}>0.99$ across all $5\times 2\times 20=200$ data points, even when $F_{\mathrm{before}}$ drops to 0.066 (Regev, $\gamma=5.0$) or 0.065 (Regev, $p=0.95$).
The observed pre-correction decay is consistent with Eq. (23): Regev ($d=64$, $1/d=0.016$) decays most steeply, while qDRIFT ($d=8$, $1/d=0.125$) decays more slowly due to non-uniform population distribution. Post-CQEC fidelity remains flat at $F\approx 1.0$ regardless of $d$ or noise strength. At $p=1$ (complete depolarizing), $\rho_{\mathrm{noisy}}=I/d$ and all off-diagonal elements vanish, so $\mathcal{D}(\rho_{\mathrm{noisy}})=\emptyset$ and recovery is impossible; the sweep stops at $p=0.95$ where modes are still preserved ($|\rho_{ij}|\geq 0.05\cdot|\rho_{ij}^{(0)}|$).
Figure 6: CQEC recovery fidelity vs. noise strength. (a) Dephasing, $\gamma\in[0.1,5.0]$. (b) Depolarizing, $p\in[0.01,0.95]$. Dashed lines: pre-correction; solid lines: post-CQEC.
VI.4 Comparison with conventional QEC
Figure 7 compares CQEC against three conventional codes across depolarizing noise $p\in[0.001,0.3]$.
Table 8: Fidelity comparison at representative error rates (CF-QPE, 4 qubits). CQEC values are asymptotic ($n\to\infty$); QEC values assume a single correction round with perfect syndrome extraction. Both represent idealized performance under their respective operational assumptions.
$p$ None Steane Surf. d=3d\!=\!3 Surf. d=5d\!=\!5 CQEC
0.01 0.989 1.000 0.994 0.998 1.000
0.10 0.905 0.987 0.979 0.974 1.000
0.20 0.811 0.949 0.918 0.848 1.000
0.30 0.718 0.885 0.824 0.639 1.000
Key observations: (1) At $p=0.01$ (below threshold), all codes perform well. (2) At $p=0.1$ (above threshold for surface codes), surface code $d=5$ degrades to 0.974 while CQEC remains at 1.000. (3) At $p=0.3$, Steane falls to 0.885, surface $d=5$ to 0.639, while CQEC maintains 1.000. The pattern is consistent across all four algorithms (Fig. 7a–d).
Figure 7: Conventional QEC vs. CQEC across depolarizing noise. Panels (a)–(d) for qDRIFT, QKAN, CF-QPE, Regev. CQEC maintains $F\approx 1$ while conventional codes degrade monotonically.
VI.5 TTN cryptographic protocol protection
Figure 8 presents a heatmap across 42 data points (6 plaintext lengths $\times$ 7 noise levels). We observe three regimes: (i) Weak noise ($\gamma\leq 0.5$, $p\leq 0.05$): pre-correction $F>0.9$, CQEC provides marginal benefit. (ii) Moderate noise ($\gamma\sim 1.5$, $p\sim 0.15$): pre-correction $F\approx 0.5$$0.7$, CQEC restores to $F=1.000$. (iii) Extreme noise ($\gamma=3.0$, $p=0.3$): pre-correction F=0.20F=0.200.280.28, approaching $1/d=0.125$; CQEC still achieves $F=1.000$ in the asymptotic limit. The recovery is uniform across all plaintext lengths, confirming that the mode structure of TTN output states is robust: all 6 plaintext lengths produce states with identical mode support $|\mathcal{D}|=56$. CQEC can restore intermediate quantum states in prepare-and-measure cryptographic schemes where the sender knows the target state, complementing the classical noise absorption mechanism in TTN protocols [34]. The practical relevance is quantum key distribution with noisy channels: CQEC applied at the receiver side can restore quantum states before measurement, improving the raw key rate under high-loss conditions.
Figure 8: TTN crypto protocol ++ CQEC protection. (a) Fidelity before CQEC (heatmap by plaintext length and noise level). (b) Fidelity after CQEC (uniformly $F=1.0$).
VI.6 Catalyst durability
Figure 9 demonstrates catalyst reuse over 100 consecutive recovery cycles (qDRIFT, dephasing $\gamma=1.5$, depolarizing $p=0.2$). The catalyst is not re-initialized between cycles: the output $\rho_{C}^{\mathrm{out}}$ from cycle $k$ is used as input for cycle $k+1$.
  • Average recovered fidelity: $F=0.990$ (all 100 cycles identical).
  • Maximum catalyst deviation: $\delta_{c}<10^{-12}$ per cycle.
  • 128-bit precision verification: $\delta_{c}<10^{-28}$, consistent with exact catalyst preservation as predicted by Theorem 2.
Since the same optimized parameters $\boldsymbol{\theta}$ are used in every cycle and the input noise is identically distributed, cycle-to-cycle stability of $F_{\mathrm{rec}}$ is expected if the catalyst state is exactly preserved. The physical content of this test is confirming that the reduced catalyst state $\mathrm{Tr}_{S}[\tau]$ remains numerically indistinguishable from the input $c$ over 100 cycles—that is, system–catalyst correlations inherent in correlated catalysis do not accumulate into marginal deviation. A more stringent test would use cycle-dependent noise realizations or feed the recovered state into a downstream computation; we leave this to future experimental work. We caution that unlimited reuse is guaranteed only in the asymptotic limit; for finite-$n$ implementations, drift of order $O(1/\sqrt{n})$ per cycle could accumulate.
Figure 9: Catalyst durability test. (a) Recovery fidelity over 100 cycles (green: recovered, red: noisy). (b) Catalyst state deviation from original (identically zero).
VI.7 Finite-copy fidelity bounds
Table 9 reports the estimated copy number $n^{*}$ required to achieve target fidelities, from Eq. (14). We estimate $C$ from the coherence spectrum of each target state: qDRIFT ($C_{\ell_{1}}=3.2$, $|\rho_{\mathrm{min}}|=0.021$), QKAN ($C_{\ell_{1}}=2.1$, $|\rho_{\mathrm{min}}|=0.18$), CF-QPE ($C_{\ell_{1}}=8.7$, $|\rho_{\mathrm{min}}|=0.0083$), Regev ($C_{\ell_{1}}=47$, $|\rho_{\mathrm{min}}|=0.0012$).
Table 9: Estimated input copies $n^{*}$ for target fidelity under dephasing ($\gamma=2.0$).
Algorithm $C$ $n^{*}(F\!\geq\!0.99)$ $n^{*}(F\!\geq\!0.999)$
qDRIFT ($d=8$) 42 $4\times 10^{5} 4.4$$4\times 10^{7}$
QKAN ($d=4$) 8.5 $1.8\times 10^{4}$ $1.8\times 10^{6}$
CF-QPE ($d=16$) 170 $7.2\times 10^{6}$ $7.2\times 10^{8}$
Regev ($d=64$) 2700 $1.8\times 10^{9}$ $1.8\times 10^{11}$
[…]
Figure 10: Finite-copy fidelity bounds from Eq. (14). (a) Fidelity vs. $n$ at $\gamma=2.0$; triangles mark $n^{*}$ for $F=0.99$. (b) Fidelity vs. $d$ at fixed $n$; empirical fit $C\approx 0.53\,d^{2.06}$. (c) Fidelity vs. $\gamma$ at $n=10^{6}$.
VI.8 DD+Twirl pipeline benchmarks
Table 10 presents the unified resource–fidelity comparison across all catalyst preparation methods under dephasing ($\gamma=2$).
Table 10: Catalyst fidelity $F_{\mathrm{cat}}$ and CQEC recovery fidelity $F_{\mathrm{rec}}$ under dephasing ($\gamma=2$). DD+Twirl uses CPMG-8 with $n=8$ copies. Parenthetical numbers after $F_{\mathrm{cat}}$ denote the copy count. Distillation $n^{*}$ is the Shiraishi–Takagi estimate for $F_{\mathrm{cat}}\geq 0.99$.
Variational ($n=0$) Standard ($n_{\mathrm{max}}$) Covariant ($n_{\mathrm{max}}$) DD+Twirl ($n=8$) Distillation
Algorithm $d$ $F_{\mathrm{cat}}$ $F_{\mathrm{rec}}$ $F_{\mathrm{cat}}$ $F_{\mathrm{rec}}$ $F_{\mathrm{cat}}$ $F_{\mathrm{rec}}$ $F_{\mathrm{cat}}$ $F_{\mathrm{rec}}$ $n^{*}$
QKAN 4 1.000 0.83 0.438 (64) 0.579 0.411 (64) 0.573 0.964 0.811 $7.3\times 10^{5}$
qDRIFT 8 0.993 0.73 0.195 (64) 0.757 0.274 (64) 0.759 0.963 0.900 $1.8\times 10^{7}$
CF-QPE 16 0.960 0.65 0.088 (32) 0.146 0.106 (32) 0.147 0.963 0.649 $2.9\times 10^{8}$
Regev 64 0.021 (8) 0.123 0.022 (8) 0.123 0.963 0.636 $7.3\times 10^{10}$
The gap between $F_{\mathrm{cat}}$ and $F_{\mathrm{rec}}$ arises from the variational recovery circuit (Sec. III.4), not the catalyst quality. With a perfect catalyst ($F_{\mathrm{cat}}=1$), the 5-gate circuit achieves $F_{\mathrm{rec}}=0.83$$1.00$ depending on $d$ (Table 7); imperfect catalysts ($F_{\mathrm{cat}}=0.96$) further reduce $F_{\mathrm{rec}}$ by $\sim\!0.1$$0.2$. In the full constructive protocol (Theorem 2), $F_{\mathrm{cat}}\to 1$ implies $F_{\mathrm{rec}}\to 1$; the variational bottleneck is the circuit expressibility, not a fundamental limit. A deeper ansatz ($\geq 15$ gates) recovers $F_{\mathrm{rec}}>0.99$ for all $d\leq 16$ even with $F_{\mathrm{cat}}=0.96$.
Key observations:
  1. 1.
    Under strong dephasing, standard and covariant swap tests alone fail ($F_{\mathrm{cat}}<0.44$). At $d=64$, the covariant method reaches only $F_{\mathrm{cat}}=0.022$—barely above $1/d$.
  2. 2.
[…]
  3. 3.
    The DD+Twirl pipeline solves the dephasing problem: $F_{\mathrm{cat}}>0.96$ uniformly across d=4d=46464 with only 8 copies—a $10^{9}$-fold improvement over distillation.
  4. 4.
    Under depolarizing noise, both standard and covariant methods achieve $F_{\mathrm{cat}}>0.93$ with 8–64 copies.
  5. 5.
    At $d=4$, the variational approach ($F_{\mathrm{cat}}=1.000$) remains competitive.
VI.9 Scaling analysis: fidelity vs. error, qubits, and copies
We analyze how catalyst fidelity scales with three key parameters under the DD+Twirl pipeline.
Fidelity vs. error strength (Fig. 11). Under raw dephasing, $F_{\mathrm{cat}}$ decays rapidly and is strongly dimension-dependent ($d=16$ reaches $F=0.1$ at $\gamma=1$). With DD+Twirl, $F_{\mathrm{cat}}$ remains above 0.9 for $\gamma\leq 2.5$ across all tested dimensions, and the curves nearly overlap—the pipeline eliminates the dimension dependence. At $\gamma=4$, DD+Twirl still maintains $F_{\mathrm{cat}}\approx 0.85$.
Figure 11: $F_{\mathrm{cat}}$ vs. dephasing strength $\gamma$ at $n=8$ copies. (a) Raw dephasing: rapid decay, strongly $d$-dependent. (b) DD+Twirl (CPMG-8): $F_{\mathrm{cat}}>0.9$ for all $d$ up to $\gamma\approx 3$, near-complete dimension independence.
Fidelity vs. system size (Fig. 12). Under raw dephasing, $F_{\mathrm{cat}}$ drops from $\sim 0.7$ at $n_{q}=1$ to $\sim 0.04$ at $n_{q}=5$ regardless of copy count. With DD+Twirl, $F_{\mathrm{cat}}$ is nearly constant: 0.97 ($n_{q}=1$), 0.96 ($n_{q}=2$55) at $n=8$. This dimension independence is a key practical advantage: the same pipeline works for any $d$ up to at least 32 without parameter tuning.
Figure 12: $F_{\mathrm{cat}}$ vs. qubit count at dephasing $\gamma=2$. (a) Raw: monotonic degradation, $F<0.05$ for $n_{q}\geq 5$. (b) DD+Twirl (CPMG-8): nearly flat at $F\approx 0.96$$0.97$, effective dimension independence.
Fidelity vs. copy count (Fig. 13). Under raw dephasing, convergence is logarithmically slow: $F_{\mathrm{cat}}\approx 0.44$ at $n=64$ for $d=4$. With DD+Twirl, the swap test recovers doubly exponential convergence: $F>0.96$ at $n=8$, $F>0.99$ at $n=32$ for all dd—approaching the distillation limit with $10^{4}$$10^{8}$ fewer copies.
Figure 13: $F_{\mathrm{cat}}$ vs. copy count at dephasing $\gamma=2$. (a) Raw: logarithmic improvement, $F<0.5$ even at $n=64$. (b) DD+Twirl (CPMG-8): doubly exponential convergence restored; $F>0.99$ at $n=32$ for all $d$.
VI.10 Resource overhead
Table 11: Resource comparison for $n_{\mathrm{logical}}$ logical qubits. CQEC qubit count is $n_{S}+n_{C}+n_{A}+n_{\mathrm{copies}}$; gate count is for the variational recovery circuit only and does not include the cost of preparing $n^{*}$ noisy copies (Table 9). With DD+Twirl, $n^{*}=8$3232, adding at most $32\times G_{\mathrm{alg}}$ gates.
$n_{\mathrm{logical}}$ CQEC (qb / gates) Steane (qb / gates) Surf. d=3d\!=\!3 (qb / gates) Surf. d=5d\!=\!5 (qb / gates)
1 8 / 5 7 / 10 13 / 36 41 / 100
3 22 / 11 21 / 30 39 / 108 123 / 300
6 137 / 20 42 / 60 78 / 216 246 / 600
The total CQEC resource cost includes: (i) preparation of $n$ noisy copies ($n\times G_{\mathrm{alg}}$ gates), (ii) DD pulses ($n\times N$ single-qubit gates), (iii) Clifford twirling ($n\times O(n_{q}^{2})$ gates), (iv) $\log_{2}n$ rounds of swap tests ($O(n_{q}\log n)$ two-qubit gates), and (v) the recovery circuit ($G_{\mathrm{CQEC}}=5$1515 EC gates). The total gate count is $n\times(G_{\mathrm{alg}}+N+O(n_{q}^{2}))+O(n_{q}\log n)+G_{\mathrm{CQEC}}$. For qDRIFT with 80 gates and $n^{*}=4.4\times 10^{5}$, the total is $\sim 3.5\times 10^{7}$ gates—more than surface code overhead. CQEC’s advantage is therefore not in absolute resource cost but in threshold-free recovery. For systems relevant to quantum advantage ($n\geq 50$), the CQEC copy overhead scales as $\mu\sim 2^{2n}e^{2\gamma}$, which is exponential—far exceeding the polynomial overhead of surface codes. A practically relevant regime for CQEC may be small quantum modules ($n\leq 10$) within a larger fault-tolerant architecture.
VII Discussion
VII.1 Two paradigms for quantum information protection
Our results reveal a complementary relationship between two protection strategies:
  1. 1.
    Classical noise absorption (as in TTN protocols [34]): The classical regression layer absorbs quantum measurement noise statistically. This requires no quantum overhead but is limited to cases where decoherence does not fundamentally alter the feature space. Under the stronger decoherence tested here ($p=0.3$, $\gamma=3.0$), pre-correction fidelity drops to $F\approx 0.20$$0.28$, which would severely degrade the quantum feature space.
  2. 2.
    Catalytic coherence amplification (CQEC): Operates directly on the quantum state before measurement. Effective for arbitrarily strong decoherence provided $C_{\ell_{1}}>0$, but requires copy overhead. CQEC restores $F=1.0$ at the state level, preserving the quantum advantage of feature extraction.
These paradigms address different noise layers: classical absorption handles measurement-level noise, while CQEC addresses state-level decoherence that classical methods cannot correct.
VII.2 Complementarity of catalyst preparation methods
The four methods occupy distinct regions of the resource–fidelity landscape:
  • Variational: best when copies are unavailable (NISQ regime, $d\leq 16$). Achieves $C_{\ell_{1}}=d-1$ (maximum coherence) but recovery fidelity is limited by the simplified recovery model.
  • Standard swap test: effective under depolarizing noise with ample copies. Achieves $F_{\mathrm{cat}}>0.93$ with $\sim\!64$ copies. Fails under dephasing.
[…]
  • DD+Twirl+Swap Test: the most powerful method under dephasing. Achieves $F_{\mathrm{cat}}>0.96$ for all tested dimensions with only 8 copies and 8 DD pulses, with negligible overhead compared to the copy-count savings.
A hybrid strategy is natural: use variational for initial construction when no copies are available, then refine via DD+Twirl as noisy copies become available. The DD+Twirl pipeline is the recommended approach for dephasing-dominated noise at any dimension.
VII.3 Limitations
Oracle access to the target state. CQEC assumes knowledge of $\rho_{0}$ for both mode verification and recovery. For Hamiltonian simulation (qDRIFT), the target is computable classically for small systems—the goal is to recover the state cheaply from noisy copies rather than re-running the expensive circuit. For factoring (Regev), knowing $\rho_{0}$ implies knowing the factorization, making CQEC circular; we include Regev solely as a stress test for high-dimensional coherence ($d=64$). Appropriate use cases include VQE states, QAOA outputs, and quantum state preparation where targets are efficiently specifiable but expensive to prepare with high fidelity.
Asymptotic nature. Theorems 1 and 2 are asymptotic: exact conversion holds only for $n\to\infty$. For finite $n$, the gap scales as $O(1/\sqrt{n})$ [31]. Our asymptotic simulations validate the existence of recovery; the finite-$n$ bounds (Table 9) quantify the practical cost, which can be prohibitive for large systems.
Comparison with conventional QEC. CQEC and QEC operate under fundamentally different assumptions. Conventional codes require only syndrome measurements without knowing the encoded state; CQEC requires the target state and $O(\mu\cdot N^{k})$ copies. A fair comparison must account for these different operational settings. The comparison in Sec. VI.4 illustrates qualitative differences, not strict superiority.
Comparison with tomography. Tomography followed by re-preparation requires $O(d^{2}/\epsilon^{2})$ measurements but destroys entanglement. CQEC preserves both coherence and entanglement structure: the covariant recovery channel acts on the original Hilbert space without collapsing the state, which is essential when the recovered state participates in subsequent entangling operations or multi-party protocols.
Gate noise in purification. Our simulations assume ideal gates. With per-gate error $\epsilon_{g}$ and $O(d\log d)$ two-qubit gates per round, the per-round circuit error is $O(\epsilon_{g}d\log d)$. For $d=64$ with current gate fidelities ($\epsilon_{g}\sim 10^{-3}$), this yields $\sim\!0.38$, indicating near-term implementation requires improved gate fidelities. For smaller systems ($d\leq 16$), superconducting platforms with $\sim\!99.5\%$ two-qubit fidelity and <100<\!100 gates per round would be viable.
Noise model. The comparison uses an effective per-qubit error model that does not capture correlations in physical noise. Correlated noise could be either favorable (preserving modes globally) or unfavorable (selectively destroying specific modes) for CQEC. Non-Markovian noise channels with memory can create time-dependent mode structures where $\mathcal{D}(\rho_{\mathrm{noisy}}(t))$ varies; CQEC would require mode verification at the time of recovery. We note that amplitude damping (included in our combined noise model) is not covariant, demonstrating that CQEC tolerates non-covariant noise channels—the covariance requirement applies only to the recovery operation $\Lambda$, not the noise.
Scalability. Our benchmarks cover d=4d=46464 (2266 qubits). For systems relevant to quantum advantage ($n_{q}\geq 50$, $d=2^{50}$), the copy overhead scales as $n^{*}\sim 2^{4n_{q}}e^{2\gamma}$, which is doubly exponential in qubit count. CQEC in its current form is therefore not a replacement for conventional QEC at scale, but rather a complementary tool for small quantum modules ($n_{q}\leq 10$) within larger architectures.
Noise model misspecification. CQEC’s mode verification (Step 1) depends on the noisy state $\rho_{\mathrm{noisy}}$, not on the noise model. If the actual noise differs from the assumed model but preserves the same modes, recovery still succeeds—the protocol is model-agnostic in this sense. However, the DD+Twirl pipeline assumes Markovian dephasing for the DD stage; under non-Markovian noise, the effective $\gamma_{\mathrm{eff}}$ may differ from $\gamma/(N+1)$, and the twirling stage still isotropizes whatever residual noise remains.
VII.4 Mode no-broadcasting and fundamental limits
Theorem 3 of Ref. [31] establishes that mode no-broadcasting prevents the creation of states with irrational coherent modes from states with only rational modes, even catalytically. If a decoherence channel selectively destroys modes with a specific energy difference $\Delta$ while preserving others, and the target state requires $\Delta$, recovery is impossible regardless of residual coherence.
We note that coherence and modes are defined with respect to the energy eigenbasis of $H$; choosing a different Hamiltonian changes which states are “coherent” and which transformations are “covariant.” In our benchmarks, $H=\sum_{q}Z_{q}$ (computational basis = energy eigenbasis), natural for qubit systems with distinct transition frequencies. This Hamiltonian has degenerate energy levels: for $n_{q}$ qubits, energy $E_{k}=n_{q}-2k$ has degeneracy $\binom{n_{q}}{k}$, creating a rich sector structure that the covariant swap test exploits (Sec. IV.3).
A physically relevant example is frequency-selective dephasing in superconducting qubits, where flux noise at frequency $\omega_{0}$ destroys coherence between levels separated by $\Delta=\hbar\omega_{0}$. This is a topological constraint on the mode lattice structure, fundamentally distinct from the quantitative error thresholds of conventional QEC.
VIII Conclusion
We have presented Catalytic Quantum Error Correction (CQEC), a quantum state recovery protocol based on the resource theory of unspeakable coherence, together with efficient catalyst preparation methods that overcome the primary resource bottleneck.
The CQEC protocol exploits the Shiraishi–Takagi result [31] that coherence can be amplified arbitrarily via catalytic covariant operations whenever the mode inclusion condition $\mathcal{C}(\rho_{0})\subseteq\mathcal{C}(\rho_{\mathrm{noisy}})$ is satisfied, with the catalyst’s reduced state preserved exactly. We demonstrated recovery across four quantum algorithms and a cryptographic protocol under three decoherence models, achieving $F>0.999$ in all 200 tested configurations in the asymptotic limit ($n\to\infty$, i.e., assuming unlimited noisy copies for the constructive protocol of Ref. [31]), and confirming the infinitely sharp mode-inclusion threshold across 10 orders of magnitude. Comparison with Steane and surface codes showed that CQEC operates without an error magnitude threshold, maintaining high fidelity at error rates where conventional codes degrade.
The main practical contribution is the DD+Twirl+Swap Test pipeline for catalyst preparation. CPMG dynamical decoupling reduces effective dephasing ($\gamma_{\mathrm{eff}}=\gamma/(N+1)$), Clifford twirling converts residual anisotropic dephasing into isotropic depolarizing noise, and the standard swap test achieves doubly exponential purification. With CPMG-8 and 8 copies, the pipeline achieves $F_{\mathrm{cat}}>0.96$ uniformly across d=4d=46464 under strong dephasing ($\gamma=2$)—a $10^{9}$-fold improvement over distillation. The total overhead is 8 DD pulses per copy plus Clifford twirling gates, negligible compared to the copy-count savings.
Our results suggest that coherence resource theory offers a promising direction for quantum state recovery that is fundamentally distinct from the stabilizer formalism, though substantial challenges remain. Three concrete next steps toward practical implementation are: (i) experimental demonstration of the DD+Twirl pipeline on superconducting or trapped-ion platforms for $d\leq 16$, where the copy overhead (n=8n=83232) is within reach; (ii) proving tight analytical bounds on the pipeline’s fidelity as a function of DD pulse count $N$, twirl sample count $K$, and copy number $n$, closing the gap between our numerical observations and rigorous guarantees; and (iii) integration with fault-tolerant architectures, where surface codes protect data qubits while CQEC restores coherence in small ancilla registers ($d\leq 32$) used for state preparation subroutines.
Acknowledgements.
Numerical simulations were performed using Qulacs [Suzuki2021qulacs] and NumPy/SciPy [numpy2020, scipy2020].
Data availability.
All simulation code (Python), numerical data (results_prxq.json, 113 KB), and plotting scripts used to generate all figures and tables are available from the corresponding author upon reasonable request, and will be deposited in a public repository upon publication. All results are deterministically reproducible with fixed random seeds.